OpenStudy (anonymous):

Represent the complex number 3 − i graphically and write it in trigonometric form.

jimthompson5910 (jim_thompson5910):

To graph 3 - i, simply graph the point (3,-1) but do so on the complex plane. Remember the complex plane is a cartesian plane with the x axis as the real number line and the y axis as the imaginary number line ------------------------------------------------------- To convert to trignometric form r*( cos(theta) + i*sin(theta) ), use the formulas theta = arctan(y/x) and r = sqrt(x^2 + y^2) --------------------------- theta = arctan(y/x) theta = arctan(-1/3) theta = -18.4349 r = sqrt(x^2 + y^2) r = sqrt(1^2 + (-3)^2) r = sqrt(10) So the number in trigonometric form is sqrt(10)*( cos(-18.4349) + i*sin(-18.4349) )

OpenStudy (anonymous):

√10(cos 208◦+ i sin 208◦) ?

OpenStudy (anonymous):

√10(cos 342◦+ i sin 342◦) ??

OpenStudy (anonymous):

√10(cos 137◦+ i sin 137◦) ???

jimthompson5910 (jim_thompson5910):

Well remember that cos(theta) = cos(theta+360) This means that there are infinitely many choices for theta But we're in luck since we only have 4 choices Notice that -18.4349 + 360 = 341.5651 which rounds to 342 (basically 342 is very close to the coterminal of -18.4349) So the answer is √10(cos 342◦+ i sin 342◦)